OpenStudy (anonymous):

Simplify. Assume that all letters denote positive numbers. \[(-3a^{1/4})(9a)^{-3/2}\]

OpenStudy (anonymous):

-3a^(1/4)3^-3a^-(3/2) =-3^-2(a)^(-5/4)

OpenStudy (anonymous):

(9a)^-3/2=3^-3a^-3/2

OpenStudy (anonymous):

=-3^-2(a)^(-5/4)

OpenStudy (angela210793):

\[-3\sqrt[4]{a}*\sqrt{(9a)^-3}=-3\sqrt[4]{a}*\sqrt{1/(9^3a^3)}\]

OpenStudy (angela210793):

Guys I'm confused...Where did -5/4 come from? oO

OpenStudy (anonymous):

1/4-3/2

OpenStudy (anonymous):

angela.....don't hurry ............u can do it........

OpenStudy (angela210793):

oh u have expressed 9 as 3^2 and then u have multiplied and then simplified right?

OpenStudy (anonymous):

yea........

OpenStudy (angela210793):

Kk Thnx ^_^ :)

OpenStudy (anonymous):

\[(-3a ^{1/4})(9a)^{-3/2}=(-3a ^{1/4})(3^{2} a)^{-3/2}= (-3a ^{1/4})3^{-3}(a)^{-3/2}\] \[=-3/81(a ^{1/4})(a)^{-3/2}\] \[=-1/9(a ^{1/4-3/2})\] \[=-1/9(a ^{-5/4})\]

OpenStudy (anonymous):

\[{-1 \over 9a^{5/4}}\] is the answer

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

\[=-1/9a ^{5/4}\]

OpenStudy (anonymous):

exactly!,u expresssed in denominator we've put it in numerator...with a negative power :)

OpenStudy (anonymous):

how do you put the straight line instead of / for divided by sign in the equation toolbox?